OpenStudy (anonymous):

What is the area of a regular hexagon with side length of 40cm? A 1200 Square root of 3. centimeters squared B 2400 Square root of 3. centimeters squared C 4800 Square root of 3. centimeters squared D 2000 Square root of 3. centimeters squared

OpenStudy (anonymous):



OpenStudy (anonymous):

The regular hexagon is made of 6 equilateral triangles. Each central angle is : 360/6 = 60 deg The side length is 40, then the side of the height OH is :40*(V3/2) = 20V3 Area of one triangle = 40* (20V3)/2 = 400*V3 Area of hexagon= 6 triangles = 6 x 400*V3 = ........